OpenStudy (anonymous):

If f (x) = x2 + 2x + 5 and g(x) = x - 1, then f (g(x)) is [1] x2 + 3x + 4 [2] x2 + 4 [3] x2 + 4x - 1 [4] x2 - 1 show work please If f (x) = x2 + 2x + 5 and g(x) = x - 1, then f (g(x)) is [1] x2 + 3x + 4 [2] x2 + 4 [3] x2 + 4x - 1 [4] x2 - 1 show work please @Mathematics

OpenStudy (anonymous):

The answer is [2] x^2 + 4 It means g(x) goes into f(x)'s x values; f(g(x)) = (x-1)^2 + 2(x-1) +5 = x^2 - 2x +1 + 2x -2 + 5 = x^2 + 4